# Pilot Contamination-Aware Graph Attention Network for Power Control in CFmMIMO

### v1

@@ III-B The Proposed Network Model
1) Preprocessing Stage: To accommodate a varying number of UEs, the inputs $\mathbf{B}$ and $\bm{\Phi}$ are first padded. Specifically, zero-padding is applied to generate $\mathbf{B}_{\rm padded}$ of size $M\times K_{\rm max}$ and $\bm{\Phi}_{\rm padded}$ of size $K_{\rm max}\times K_{\rm max}$. For simplicity of notation, $\mathbf{B}$ and $\bm{\Phi}$ are used to refer to their padded counterparts, and $K=K_{\rm max}$ in the following discussion unless explicitly stated.
Given that the values of large-scale fading coefficients span different orders of magnitude, the natural logarithm operation is applied to each element of $\mathbf{B}$, resulting in $\mathbf{B}_{\rm log}$. Following this, $\mathbf{B}_{\rm log}$ is transformed to have zero mean and unit variance across all elements, resulting in $\mathbf{B}_{\rm std}$. Next, scaling and shifting are performed on $\mathbf{B}_{\rm std}$ along each column, which can be expressed as
$\mathbf{B}_{\rm pre}=\left(\bm{\alpha}\mathbf{1}_{K}^{\rm T}\right)*\mathbf{B}_{\rm std}+\bm{\beta}\mathbf{1}_{K}^{\rm T},$ (12)
@@
2) Attention Mechanism Block: The attention mechanism allows the model to focus on specific parts of the input data by assigning varying importance to different components. To capture the interrelationship among nodes, a masked single-head attention mechanism is introduced, where masking enables the network to adapt to scenarios where the number of UEs varies. The attention mechanism block comprises $T=4$ hidden layers, with $\mathbf{x}_{i}(0)$ as the input to the first attention layer, the output sizes of $4$ layers are $32,64,64,64$, respectively. A detailed explanation of the attention mechanism block is provided in Subsection III-C.
@@
Next, the node representation vectors undergo another linear transformation to map them into values. Following this, additional processing steps–similar to those in [7]–are applied, including the operations described in (14) and (15), giving
@@
where ${\rm SoftPlus}(\cdot)$ denotes the Softplus activation function, and ${\mathcal{L}}_{3}$ is a linear function that maps a vector to a scalar value. To extract the power control coefficients for the active users while ensuring that the results satisfy the constraints in (5), masking and projection operation ${\rm Proj}_{\mathcal{S}}(\cdot)$ in [13] are applied, yielding the power control coefficients as
$\hat{\mathbf{M}}={\rm Proj}_{\mathcal{S}}({\mathbf{Y}}_{2}\bm{\Phi}_{\rm mask}),$ (15)
where $\mathbf{Y}_{2}$ is obtained by reshaping the vector $\mathbf{y}_{2}$ into a matrix of size $M\times K$, and the diagonal matrix $\bm{\Phi}_{\rm mask}$ shares the same diagonal elements as $\bm{\Phi}$.
@@ III-C Attention Mechanism for Graph Neural Network
$\mathbf{x}_{i}(t)=\mathrm{Norm}(\mathrm{ReLU}(\mathbf{y}_{{\rm AP},i}(t)+\mathbf{y}_{{\rm UE},i}(t))),i=1,\ldots,MK,$ (16)
where $\mathrm{Norm}(\cdot)$ denotes the normalization operation. Here $\mathbf{y}_{{\rm AP},i}(t)$ represents the component of the update through the attention mechanism, based on the information of nodes that are associated with the same APs as the $i$th node, and it can be expressed as
$\mathbf{y}_{{\rm AP},i}(t)=\mathcal{L}^{{\rm AP},t}_{1}\left(\mathbf{x}_{i}(t-1)\right)+$ (17)
$\mathcal{L}^{{\rm AP},t}_{5}\left(\sum\limits_{j\in\bm{\mathcal{V}}_{\rm AP},i}\psi_{ij}\alpha_{{\rm AP},ij}\left[\mathcal{L}^{{\rm AP},t}_{2}\left(\mathbf{x}_{j}(t-1)\right)+\mathcal{L}_{{\psi}}^{t}\left(\psi_{ij}\right)\right]\right),$
$\alpha_{{\rm AP},ij}=$ (18)
[…]
where $\mathcal{L}^{{\rm AP},t}_{q}$ ($q=1,\cdots,5$) denotes $5$ different linear transformations for capturing the information from AP-type edges in the $t$th layer, while $\mathcal{L}_{{\psi}}^{t}$ is a linear transformation used to capture pilot contamination information in the $t$th layer. Moreover, ${\bm{\mathcal{V}}}_{{\rm AP},i}$ denotes the set of nodes that share the same AP as the $i$th node, $\mathbf{a},\mathbf{b}\right>=\exp({\mathbf{a}^{\rm T}\mathbf{b}}/{\sqrt{D}})$, and $D$ is the length of vector $\mathbf{a}$.
Note that $\mathbf{y}_{{\rm UE},i}(t)$ is obtained through a similar process, but without considering pilot contamination. Specifically, $\mathcal{L}_{{\psi},t}\left(\cdot\right)$ is removed, as pilot contamination only arises between users.
@@ IV-A Simulation Setup
In this simulation, the large-scale fading coefficient $\beta_{mk}$ in $\rm dB$ is modeled as
$\beta_{mk}=\mathrm{PL}_{mk}+z_{mk},$ (19)
where $\mathrm{PL}_{mk}$ denotes the path loss in $\rm dB$, modeled using a three-slope model, and $z_{mk}$ accounts for shadow fading, following $\mathcal{CN}(0,\sigma_{\rm sh}^{2})$. The same large-scale fading parameters as in [2] are used to characterize the propagation path. The noise signal is simulated using the same parameters as in [7].
@@
IV-B Methods Compared and Performance Metrics
We compared the performance of the proposed method with two other power allocation methods: the APG method and the transformer-based method [^2] The proposed network and transformer-based network were trained on the NVIDIA H200 GPU.. The APG method, a traditional optimization-based method as described in [13], serves as a baseline to evaluate the performance of the proposed graph attention network, with the smoothness parameter $\lambda=3$ in (11). The transformer-based method of [7], a learning-based method also referred to as PAPC, is included for comparison as well. Considering that all existing GNN-based methods thus far are under the assumption of free pilot contamination, whereas our proposed method highlights its ability to handle pilot contamination, the GNN method is not included for comparison here.
The empirical cumulative distribution function (CDF) of the per-UE SE serves as a tool to visualize the SE distribution. In the context of the max-min fairness objective, a steeper CDF curve that shifts towards the right indicates a more desirable result, as it suggests improved fairness across UEs, as well as higher SEs for all UEs. Besides, we provide the running times of methods to assess their feasibility in real-time applications.
IV-C Performance Comparison in Different Scenarios
@@
TABLE I: Parameters for $5$ Scenarios
@@
Pilot Contamination $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Varying UEs $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Training Samples 50K 100K 800K 100K 800K
The CDF results for $5$ scenarios are presented in Fig. 2. Note that the proposed Graph ATtention network method is labeled as GAT in Figs. 2 and 3 for simplicity. It can be observed that the curve of the proposed method is to the right of those of both APG and PAPC in all scenarios, demonstrating its adaptability across different application scenarios, including variations in communication network size, pilot contamination, and whether the number of UEs is fixed or not. It is noteworthy that in Scenarios 3 and 5, the curves do not ascend as sharply as those of the other three scenarios. This is because the number of active UEs varies for each test sample, leading to a wider range of SE per UE.
@@
Table II presents the runtime for each method across different scenarios. The CPU measurements are obtained on a $64$-bit Windows system with $16$ GB RAM and an Intel Core i5 ($2.60$ GHz), while the GPU measurements are obtained using an NVIDIA A40. As shown, compared to $7.938\,[{\rm s}]$ processing time for APG in a scenario with $64$ APs and $40$ UEs, the proposed method achieves a processing time of under $46\,[{\rm ms}]$ on the GPU despite $552\,[{\rm ms}]$ on the CPU, highlighting the real-time performance of the learning-based method and demonstrating its potential for power control in real-time applications.
TABLE II: Runtime comparison between different methods (in seconds).
@@
APG on CPU $1.667$ $3.032$ $7.938$
Proposed on CPU $0.009$ $0.063$ $0.552$
Proposed on GPU $0.003$ $0.004$ $0.046$
Figure 2: Result comparison between $3$ methods across $5$ scenarios.

### v4

@@ III-B The Proposed Network Model
1) Preprocessing Stage: Given that the values of large-scale fading coefficients span different orders of magnitude, the natural logarithm operation is applied to each element of $\mathbf{B}$, resulting in $\mathbf{B}^{\rm log}$. Following this, $\mathbf{B}^{\rm log}$ is transformed to have zero mean and unit variance across all elements, resulting in $\mathbf{B}^{\rm std}$. Next, scaling and shifting are performed on $\mathbf{B}^{\rm std}$ along each column, which can be expressed as
$\mathbf{B}^{\rm pre}=\left(\bm{\alpha}\mathbf{1}_{K}^{\rm T}\right)*\mathbf{B}^{\rm std}+\bm{\beta}\mathbf{1}_{K}^{\rm T},$ (12)
@@
2) Attention Mechanism Block: The attention mechanism allows the model to focus on specific parts of the input data by assigning varying importance to different components. To capture the interrelationship among nodes, a single-head attention mechanism is introduced. The attention mechanism block comprises $T=4$ hidden layers, with $\mathbf{x}_{i}(0)$ as the input to the first attention layer, the output sizes of $4$ layers are $32,64,64,64$, respectively. A detailed explanation of the attention mechanism block is provided in Subsection III-C.
@@
Next, the node representation vectors undergo another linear transformation to map them into values. Following this, additional processing steps—similar to those in [7]—are applied, including the operations described in (14) and (15), giving
@@
where ${\rm SoftPlus}(\cdot)$ denotes the Softplus activation function, and ${\mathcal{L}}_{3}$ is a linear function that maps a vector to a scalar value. To ensure that the results satisfy the constraints in (5), projection operation ${\rm Proj}_{\mathcal{S}}(\cdot)$ in [14] are applied, yielding the power control coefficients as
$\hat{\mathbf{M}}={\rm Proj}_{\mathcal{S}}({\mathbf{Y}}_{2}),$ (15)
where $\mathbf{Y}_{2}$ is obtained by reshaping the vector $\mathbf{y}_{2}$ into a matrix of size $M\times K$.
@@ III-C Attention Mechanism for Graph Neural Network
$\mathbf{x}_{i}(t)=\mathrm{Norm}(\mathrm{ReLU}(\mathbf{y}_{i}^{\rm AP}(t)+\mathbf{y}_{i}^{\rm UE}(t))),$ (16)
where $\mathrm{Norm}(\cdot)$ denotes the normalization operation. Here $\mathbf{y}_{i}^{\rm AP}(t)$ represents the component of the update through the attention mechanism, based on the information of nodes that are associated with the same APs as the $i$th node, and it can be expressed as
$\mathbf{y}_{i}^{\rm AP}(t)=\mathcal{L}^{{\rm AP},t}_{1}\left(\mathbf{x}_{i}(t-1)\right)+$ (17)
$\mathcal{L}^{{\rm AP},t}_{5}\left(\sum\limits_{j\in\bm{\mathcal{V}}_{i}^{\rm AP}}\alpha_{i,j}^{\rm AP}\left[\mathcal{L}^{{\rm AP},t}_{2}\left(\mathbf{x}_{j}(t-1)\right)+\mathcal{L}_{{\varphi}}^{t}\left(\varphi_{i,j}\right)\right]\right),$
$\alpha_{i,j}^{\rm AP}=$ (18)
$\frac{\left<{\mathcal{L}^{{\rm AP},t}_{3}\left(\mathbf{x}_{i}(t-1)\right)},\mathcal{L}^{{\rm AP},t}_{4}\left(\mathbf{x}_{j}(t-1)\right)+\mathcal{L}_{{\varphi}}^{t}\left(\varphi_{i,j}\right)\right>}{\sum\limits_{u\in{\bm{\mathcal{V}}_{i}^{\rm AP}}}\left<{\mathcal{L}^{{\rm AP},t}_{3}\left(\mathbf{x}_{i}(t-1)\right)},\mathcal{L}^{{\rm AP},t}_{4}\left(\mathbf{x}_{u}(t-1)\right)+\mathcal{L}_{{\varphi}}^{t}\left(\varphi_{i,u}\right)\right>},$
where $\mathcal{L}^{{\rm AP},t}_{q}$ ($q=1,\cdots,5$) denotes $5$ linear transformations for capturing the information from AP-type edges in the $t$th layer, while $\mathcal{L}_{{\varphi}}^{t}$ is a linear transformation used to capture pilot contamination information in the $t$th layer. Moreover, ${\bm{\mathcal{V}}}_{i}^{\rm AP}$ denotes the set of nodes that share the same AP as the $i$th node, $\left<\mathbf{a},\mathbf{b}\right>=\exp({\mathbf{a}^{\rm T}\mathbf{b}}/{\sqrt{D}})$, and $D$ is the length of $\mathbf{a}$.
Note that $\mathbf{y}_{i}^{\rm UE}(t)$ is obtained through a similar process, but without considering pilot contamination. Specifically, $\mathcal{L}_{{\varphi}}^{t}\left(\cdot\right)$ is removed, as pilot contamination only arises between users.
@@ IV-A Simulation Setup
In this simulation, the large-scale fading coefficient $\beta_{m,k}$ in $\rm dB$ is modeled as
$\beta_{m,k}=\mathrm{PL}_{m,k}+z_{m,k},$ (19)
where $\mathrm{PL}_{m,k}$ denotes the path loss in $\rm dB$, modeled using a three-slope model, and $z_{m,k}$ accounts for shadow fading, following $\mathcal{CN}(0,\sigma_{\rm sh}^{2})$. The same large-scale fading parameters as in [2] are used to characterize the propagation path. The noise signal is simulated using the same parameters as in [7].
@@
IV-B Dataset Preparation and Padding Strategy
To address the requirement that all tensors within a batch must have the same shape in modern deep-learning frameworks such as PyTorch, which is used in our implementation, we employ a zero-padding strategy to handle training datasets with a varying number of UEs. Specifically, $K$ is set as the maximum number of UEs, $K_{\rm max}$, within the coverage area. If the actual number of UEs, $K_{\rm act}$, is smaller than $K_{\rm max}$, an extra $K_{\rm max}-K_{\rm act}$ UEs are added as padding. During training data preparation, zero-padding is applied to $\mathbf{B}$ and $\bm{\Phi}$, producing a new $\mathbf{B}$ of size $M\times K_{\rm max}$ and a new $\bm{\Phi}$ of size $K_{\rm max}\times K_{\rm max}$ as actual inputs, ensuring uniform input dimensions across training batches. Accordingly, the following modifications are made to eliminate the influence of padded UEs: in (15), ${\mathbf{Y}}_{2}$ is replaced with ${\mathbf{Y}}_{2}\bm{\Phi}_{\rm mask}$; in (17), $\alpha_{i,j}^{\rm AP}$ is substituted with $\varphi_{j,j}\alpha_{i,j}^{\rm AP}$. Here, the diagonal matrix $\bm{\Phi}_{\rm mask}$ shares the same diagonal elements as the padded $\bm{\Phi}$.
IV-C Methods Compared and Performance Metrics
We compared the performance of the proposed method with two other power allocation methods: the APG method and the transformer-based method [^2] The proposed network and transformer-based network were trained on the NVIDIA H200 GPU.. The APG method, a traditional optimization-based method as described in [14], serves as a baseline to evaluate the performance of the proposed graph attention network, with the smoothness parameter $\lambda=3$ in (11). The transformer-based method of [7], a learning-based method also referred to as PAPC, is included for comparison as well. Considering that all existing GNN-based methods thus far are under the assumption of free pilot contamination, whereas our proposed method highlights its ability to handle pilot contamination, GNNs are not included for comparison.
The empirical cumulative distribution function (CDF) of the per-UE SE serves as a tool to visualize the SE distribution. Besides, we provide the running times of methods to assess their feasibility in real-time applications.
IV-D Performance Comparison in Different Scenarios
@@
TABLE I: Parameters for 55 Scenarios
@@
Pilot Contamination
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Varying UEs
$\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Training Samples
50K 100K 800K 100K 800K
Figure 2: Result comparison between 33 methods across 55 scenarios.
The CDF results for $5$ scenarios are presented in Fig. 2, and the quantitative results corresponding to Eq.(11) is provided in Table II. These values in Table II represent the average results over all test samples. Note that the proposed Graph ATtention network method is labeled as GAT in Figs. 2 and 3 for simplicity. In the context of the max-min fairness objective, a steeper CDF curve that shifts towards the right indicates a more desirable result, as it suggests improved fairness across UEs, as well as higher SEs for all UEs. Observing Fig. 2 and Table II, it can be concluded that the proposed method outperforms both APG and PAPC in all scenarios, including variations in communication network size, pilot contamination, and whether the number of UEs is fixed or not. Note that in Scenarios 3 and 5, the curves do not ascend as sharply as those of the other three scenarios. This is because the number of active UEs varies for each test sample, leading to a wider range of SE per UE.
TABLE II: Quantitative results showing the corresponding values of ($1$) (3 methods, 55 scenarios).
Scenarios APG PAPAC Proposed
Scen. 1 1.9647 1.9641 1.9731
Scen. 2 1.7524 1.7363 1.7619
Scen. 3 2.0219 2.0132 2.0430
Scen. 4 1.6311 1.6091 1.6380
Scen. 5 1.9327 1.9314 1.9541
@@
Table III presents the runtime for each method across different scenarios. The CPU measurements are obtained on a $64$-bit Windows system with $16$ GB RAM and an Intel Core i5 ($2.60$ GHz), while the GPU measurements are obtained using an NVIDIA A40. As shown, compared to $5.5185\,[{\rm s}]$ GPU processing time for APG in Scenario 5 with $64$ APs and $40$ UEs, the proposed method achieves a processing time of $2.2\,[{\rm ms}]$ on the GPU despite $592\,[{\rm ms}]$ on the CPU, highlighting the real-time performance of the learning-based method and demonstrating its potential for power control in real-time applications. Note that although the PAPC’s is slightly less than the proposed model, the performance improvement justifies the negligible time difference.
While TDD systems leverage channel reciprocity, practical implementations can encounter errors due to hardware impairments or calibration mismatches, leading to large-scale channel estimation errors. To assess the impact of large-scale channel estimation errors on the performance of the proposed model, another experiment was conducted, with the result is presented in Fig. 4, and the experimental settings are as follows. Let $\beta_{m,k}^{[\rm dB]}=10\log_{10}(\beta_{m,k})$ denote the large-scale fading coefficient in the decibel scale. To model the effect of estimation errors in the simulations, we add additive white Gaussian noise to the channel coefficients, which allows us to realistically capture the impact of imperfect channel knowledge on system performance [21, 22, 23]. The corresponding estimated values with errors can then be modeled as
$\hat{\beta}_{m,k}^{[\rm dB]}=\beta_{m,k}^{[\rm dB]}+\mathcal{CN}(0,\sigma^{2}).$
To evaluate the impacts of estimation errors, we consider different values of $\sigma$: $0.5,1,1.5$, and $2$. The experiments were conducted under Scenario 2, and the resulting CDFs are given in Fig. 4. As shown in Fig. 4, the performance degrades with increasing $\sigma$, as reflected by the CDF results. However, the proposed model consistently achieves nearly the same performance as the APG algorithm, even when $\sigma$ is as high as 2, which validates the effectiveness and robustness of our proposed method.
Next, to assess the performance gain introduced by incorporating pilot contamination information, an ablation study is conducted by removing the pilot-related input, $\Phi$, from the proposed model. Using Scenarios 2 and 4 as representative examples, the resulting CDFs of the per-user spectral efficiency are illustrated in Fig. 5. As observed, the CDF curves of the GAT model with pilot contamination awareness (in solid line) in both scenarios are significantly shifted to the right compared to its counterpart without pilot information (in dashed line). This clearly demonstrates the importance of including pilot contamination information in the GNN input, highlighting its critical role in enhancing power control performance in CF-mMIMO systems.
In conclusion, the proposed method achieves performance comparable to APG. Moreover, it significantly reduces computational complexity, highlighting the superior suitability of GNNs for CFmMIMO applications.
TABLE III: Runtime comparison between methods (in seconds).
@@
CPU
GPU
CPU
GPU
CPU
GPU
APG 1.5626 1.1391 3.5581 2.4226 11.2485 5.5185
PAPC 0.0022 0.0012 0.0134 0.0013 0.0167 0.0014
Proposed 0.0083 0.0020 0.0613 0.0021 0.5920 0.0022
Figure 4: Performance comparison for different estimation errors in Scen. 2
Figure 5: Ablation study on pilot information awareness.